# SLIDE: Simultaneous Model Downloading and Inference at the Wireless Network Edge

### v1

@@ I Introduction
With mounting concerns about privacy and the increasing computing power of mobile devices, on-device inference has become a prevalent paradigm, enabling users to run AI models directly on their end devices [1, 2, 3]. In this respect, AI model downloading has emerged as an essential service in next-generation mobile networks to support ubiquitous on-device inference [4, 5, 6]. Unlike the traditional model (or mobile app) downloading, which is often time-consuming, the future model downloading is expected to be a real-time service [7, 8]. For instance, the 5G standard specifies model downloading latency requirements as low as 1 second for time-sensitive applications such as autonomous driving and robotics [4, 9]. Real-time model downloading benefits on-device inference in two aspects. (i) Storage efficiency: Due to the diverse range of AI applications and the large size of models [10, 11, 12], users prefer not to, or simply cannot, pre-store all the AI models that they might need in advance. For example, Google’s on-device language model Gemini Nano-2 remains 1.51 GB even after 4-bit quantization [13]. Real-time model downloading enables users to fetch models on demand from off-site locations, thereby saving on-device storage. (ii) Right-on-time intelligence: Real-time downloading is crucial for time-sensitive applications when pre-downloading is infeasible. For example, mobile network is expected to download region-specific models to users, such as localized models for augmented reality [14] and autonomous driving [15, 16], which are adapted/optimized for this area. In such cases, since the downloading of the context-aware model can occur only when users enter the region, real-time model downloading is essential to ensure low service latency for users.
@@ II Related Work
Model downloading: In existing DAI frameworks, users download the required AI models from edge servers (e.g., BSs) for local inference [26]. To reduce model downloading latency, Wu et al. leveraged parameter sharing across models to enable BS to broadcast shared model blocks to users [17]. Moreover, to better utilize computing resources of both user devices and edge servers, cooperative edge inference splits AI models into user-side and server-side sub-models [27, 28], allowing users to download and run the user-side sub-model and upload intermediate features to the server for further inference [29, 30]. In these studies, on-device inference begins only after model downloading completes, which can hence be reduced to conventional resource allocation problems. However, these approaches fall short for SLIDE since they do not consider the overlapping of model downloading and local inference. This overlapping introduces recursive dependencies across layers in the E2E latency of SLIDE, making the latency jointly influenced by bandwidth allocation, layer sizes, and computing workload. This indicates that SLIDE requires a tailored resource allocation strategy.
@@
Video streaming: Video streaming techniques divide videos into multiple segments, allowing users to download the video segments and consume the video simultaneously [31, 32, 33]. This progressive downloading enables users to start playback before downloading the full video. However, this approach is inapplicable to downloading and inference. Although AI models can similarly be divided into layers, users can only benefit from the models upon inference completion for the entire model. Moreover, unlike video segments, which are uniformly timed [34, 35], computing workloads and data sizes vary across layers. These factors prevent the application of video streaming schemes to SLIDE.
Layer-wise inference: In computer architecture, layer-wise inference sequentially transfers model layers from disk/system memory to GPU memory, allowing the GPU to perform inference with each layer once being loaded [36, 37]. Similar to SLIDE, loading of the subsequent layer can begin concurrently with the computation of the current layer [38]. However, these methods focus on on-device inference and overlook the challenges of model downloading in wireless edge networks, where users share limited spectrum. In such settings, downlink rates are typically limited to hundreds of Mbps [19], far below the several to tens of GB/s available for on-device model loading [39]. Thus, joint spectrum and computing resource allocation must be judiciously determined to support model downloading and inference in multi-user scenarios in wireless edge networks.
@@ III-A Network Scenario
Although a user can perform inference using the downloaded models multiple times, our optimization framework focuses on optimizing the first-time model downloading when the user does not store the model locally. If users pre-store the needed models, they can directly perform on-device inference without sending their requests to the BS.
@@ III-B The Procedure of SLIDE
The detailed procedure of SLIDE, as demonstrated in the lower part of Fig. 3, is illustrated as follows. Suppose that user $k$ is provisioned with model $i$, with $\left\{1,2,\dots,L_{i}\right\}$ being the ordered set of layers of model $i$. Starting from time 0, user $k$ continuously downloads the layers of model $i$ in the order of $l_{i}$ to the last layer $L_{i}$, while simultaneously performing layer-wise inference using the already downloaded layers. To further reduce the E2E latency, user $k$ moves task data to the GPU memory and executes model instantiation when downloading the first layer. Intuitively, the inference of user $k$ for layer $l_{i}$ begins at the latest of the following two events: (i) the completion of downloading layer $l_{i}$; and (ii) the completion of inference for the preceding layer $l_{i}-1$. As an exception, inference for layer 1 starts after the layer has been downloaded and the process of data movement and model instantiation is completed. Once the inference for layer $L_{i}$ is completed, user $k$ obtains the final inference result.
@@
SLIDE supports both feedforward models (e.g., convolutional neural networks (CNNs) like ResNet and Vision Transformers (ViTs)) and autoregressive models (e.g., large language models (LLMs)). For autoregressive models, SLIDE overlaps the layer downloading with the prefilling stage, in which all input tokens are processed before generating the first output token. Moreover, while SLIDE focuses on full model downloads, it can extend to cases where users update base models by downloading task-specific parameters, e.g., LoRA matrices for each LLM layer [40].
@@ III-C2 Layer inference latency
where $V_{1}\left(k,l_{i}\right)$ denotes the time required for user $k$ to transfer layer $l_{i}$ from system memory to GPU memory [^2] Without loss of generality, the GPU memory transfer time is represented by $V_{1}(k,l_{i})$, typically approximated as the ratio of the data size of layer $l_{i}$ to the data rate between the system memory and GPU memory [41].. The second term in (4) represents the time required for user $k$ to perform the forward propagation of layer $l_{i}$ on the GPU [42, 43, 44]. Moreover, $b_{k}$ denotes the inference batch size of the input data samples of user $k$, $W_{l_{i}}$ represents the computation workload (in FLOPs) required by layer $l_{i}$ to process a single data sample, $\kappa_{k}$ denotes the computing intensity (in cycles/FLOP) of user $k$, and $f_{k}$ is the GPU clock frequency (in cycles/s) of user $k$. $z_{k,l_{i}}\in\left[0,1\right]$ is the GPU frequency allocation scaling factor of user $k$ for the forward propagation of layer $l_{i}$, satisfying
@@ IV-A1 Computing resource allocation for user $k$ with model $i$ given $y_{k}$
IV-A2 Model provisioning and computing resource allocation for user $k$ under given $y_{k}$
@@ V-A Experimental Setup
  • Conventional DAI: the conventional downloading and inference approach, where users start inference only after the entire model is downloaded [53].

### v2

@@ I Introduction
With mounting concerns about privacy and the increasing computing power of mobile devices, on-device inference has become a prevalent paradigm, enabling users to run AI models directly on their end devices [1, 2]. In this respect, AI model downloading has emerged as an essential service in next-generation mobile networks to support ubiquitous on-device inference [3, 4]. Unlike the traditional model (or mobile app) downloading, which is often time-consuming, the future model downloading is expected to be a real-time service [5]. For instance, the 5G standard specifies model downloading latency requirements as low as 1 second for time-sensitive applications such as autonomous driving and robotics [3]. Real-time model downloading benefits on-device inference in the following aspects. (i) Storage efficiency: Due to the diverse range of AI applications and the large size of models, users prefer not to, or simply cannot, pre-store all the AI models that they might need in advance. For example, Google’s on-device language model Gemini Nano-2 remains 1.51 GB even after 4-bit quantization [6]. Real-time model downloading enables users to fetch models on demand from off-site locations, thereby saving on-device storage. (ii) Just-in-time intelligence: Real-time downloading is crucial for time-sensitive applications when pre-downloading is infeasible, particularly when user demands are unpredictable or specific model versions are not available in advance. For example, mobile network is expected to download region-specific models to users, such as localized models for augmented reality [7] and autonomous driving [8]. In such cases, since the downloading of the context-aware model can occur only when users enter the region, real-time model downloading is essential to ensure low service latency for users. Moreover, while recent advances in large language models (LLMs) have promoted the deployment of on-device foundation models that can be reused across multiple applications, such a paradigm does not eliminate the need for real-time model parameter downloading. In practice, since foundation models face inherent tradeoffs between model size and generality, they cannot be optimally specialized for all tasks under limited resource budgets. Therefore, users often request task-specific parameters (such as LoRA matrices [9]) for their requested tasks. As a result, dynamic and real-time model downloading is essential for providing flexible, timely, and up-to-date on-device intelligence to end users.
@@ II Related Work
Model downloading: In existing DAI frameworks, users download the required AI models from edge servers (e.g., BSs) for local inference [19]. To reduce model downloading latency, Wu et al. leveraged parameter sharing across models to enable BS to broadcast shared model blocks to users [10]. Moreover, to better utilize computing resources of both user devices and edge servers, cooperative edge inference splits AI models into user-side and server-side sub-models [20], allowing users to download and run the user-side sub-model and upload intermediate features to the server for further inference [21, 22]. In these studies, on-device inference begins only after model downloading completes, which can hence be reduced to conventional resource allocation problems. However, these approaches fall short for SLIDE since they do not consider the overlapping of model downloading and local inference. This overlapping introduces recursive dependencies across layers in the E2E latency of SLIDE, making the latency jointly influenced by bandwidth allocation, layer sizes, and computing workload. This indicates that SLIDE requires a tailored resource allocation strategy.
@@
Video streaming: Video streaming techniques divide videos into multiple segments, allowing users to download the video segments and consume the video simultaneously [23, 24]. This progressive downloading enables users to start playback before downloading the full video. However, this approach is inapplicable to downloading and inference. Although AI models can similarly be divided into layers, users can only benefit from the models upon inference completion for the entire model. Moreover, unlike video segments, which are uniformly timed [25, 26], computing workloads and data sizes vary across layers. These factors prevent the application of video streaming schemes to SLIDE.
Layer-wise inference: In computer architecture, layer-wise inference sequentially transfers model layers from disk/system memory to GPU memory, allowing the GPU to perform inference with each layer once being loaded [27, 28]. Similar to SLIDE, loading of the subsequent layer can begin concurrently with the computation of the current layer [29]. Moreover, in wireless edge networks, AI models can be partitioned into user-side and server-side sub-models for split inference by uploading intermediate features from end devices to edge servers [30]. To improve efficiency, this framework executes pipelined inference, allowing user-side inference to proceed in parallel with server-side inference on previously uploaded features. However, these methods focus on on-device inference and overlook the challenges of model downloading in wireless edge networks, where users share limited spectrum. In such settings, downlink rates are typically limited to hundreds of Mbps [12], far below the several to tens of GB/s available for on-device model loading [31]. Thus, joint spectrum and computing resource allocation must be judiciously determined to support model downloading and inference in multi-user scenarios in wireless edge networks.
@@ III-A Network Scenario
The frequently used notations are summarized in Table I.
Remark 1.
Although a user can perform inference using the downloaded models multiple times, our optimization framework focuses on optimizing the first-time model downloading when the user does not store the model locally. As discussed in the Introduction, it is generally infeasible to pre-store or persistently cache all potentially requested models. Consequently, model cache misses are unavoidable in practice, and SLIDE serves as an enabling step to efficiently deliver the missing model from the BS by overlapping model downloading with inference, after which model caching, prefetching, and reuse can take place.
@@
TABLE I: Frequently Used Notations
Symbol Description
$\mathcal{K}$ Set of users.
$\mathcal{I}$, $\mathcal{I}_{k}$ Model library and set of models requested by user $k$.
$K$, II, $L_{i}$ Number of users, models, and layers of model $i$.
$k$, $i$, $l_{i}$ User index, model index, and layer index of model $i$.
$x_{k,i}\in{\bf{X}}$ Model provisioning indicator variable.
$y_{k}\in{\bf{Y}}$ Spectrum bandwidth allocation for user $k$.
$z_{k,l_{i}}\in{\bf{Z}}$ GPU frequency allocation scaling factor of user $k$ for layer $l_{i}$.
${\bf{Z}}_{k,i}$ Vector of per-layer $z_{k,l_{i}}$ for user $k$ and model $i$.
$S_{l_{i}}$ Data size of layer $l_{i}$.
$B$ Total bandwidth of the BS.
$R_{k}$, $\gamma_{k}$ Spectral efficiency and channel gain between user $k$ and the BS.
$P$, $N_{0}$ Transmit power spectral density and noise power spectral density.
$\bar{T}_{k}$, $Q_{k}$ End-to-end latency requirement and maximum inference energy budget of user $k$.
$\tau_{k,l_{i}}$ Time for user $k$ to download layer $l_{i}$.
$T_{k,i}\left(z_{k,l_{i}}\right)$ Inference latency of user $k$ for layer $l_{i}$.
$V_{1}\left(k,l_{i}\right)$ Time required for user $k$ to transfer layer $l_{i}$ from system memory to GPU memory.
$W_{l_{i}}$ Computation workload of layer $l_{i}$ for processing one data sample.
$b_{k}$, $\kappa_{k}$, $f_{k}$, $\Psi_{k}$ Inference batch size, computing intensity, GPU clock frequency, and power coefficient of user $k$.
$t_{k,l_{i}}\left(y_{k},{\bf{Z}}_{k,i}\right)$, $t_{k,L_{i}}\left(y_{k},{\bf{Z}}_{k,i}\right)$ Time when user $k$ obtains the output of layer $l_{i}$ and the end-to-end latency of user $k$.
$e_{k,i}\left({\bf{Z}}_{k,i}\right)$ Inference energy consumption of user $k$ for model $i$.
$e_{1}\left(k,i\right)$ Energy consumption of user $k$ for model instantiation, data movement, and model transfer to GPU memory.
$e_{k}\left({\bf{X}}_{k},{\bf{Z}}_{k}\right)$ Total inference energy consumption of user $k$.
$x^{*}_{k}\in{\bf{X}}^{*}$, $y^{*}_{k}\in{\bf{Y}}^{*}$, $z^{*}_{k,l_{i}}\in{\bf{Z}}^{*}$ Optimal model provisioning, spectrum bandwidth allocation, and computing resource allocation.
@@ III-B The Procedure of SLIDE
The detailed procedure of SLIDE, as demonstrated in the lower part of Fig. 3, is illustrated as follows. Suppose that user $k$ is provisioned with model $i$, with $\left\{1,2,\dots,L_{i}\right\}$ being the ordered set of layers of model $i$. In SLIDE, a layer represents a model component (e.g., an individual model layer, a block, or a group of consecutive blocks), depending on the model architecture and chosen granularity. Starting from time 0, user $k$ continuously downloads the layers of model $i$ in the order of $l_{i}$ to the last layer $L_{i}$, while simultaneously performing layer-wise inference using the already downloaded layers. To further reduce the E2E latency, user $k$ moves task data to the GPU memory and executes model instantiation when downloading the first layer. Intuitively, the inference of user $k$ for layer $l_{i}$ begins at the latest of the following two events: (i) the completion of downloading layer $l_{i}$; and (ii) the completion of inference for the preceding layer $l_{i}-1$. As an exception, inference for layer 1 starts after the layer has been downloaded and the process of data movement and model instantiation is completed. Once the inference for layer $L_{i}$ is completed, user $k$ obtains the final inference result. Note that for model components whose internal computation is non-sequential (e.g., residual connections or parallel sub-module computation), SLIDE abstracts each such component as a single layer and incorporates its internal computation into the computation latency of that layer.
@@
SLIDE supports feedforward models (e.g., convolutional neural networks (CNNs) like ResNets and Vision Transformers (ViTs)), autoregressive models (e.g., LLMs), and recurrent neural networks (RNNs). SLIDE enables parameter downloading for subsequent layers to overlap with the inference of the current layer. For autoregressive models, this overlap occurs during the prefilling stage, where input tokens are processed in a layer-wise feedforward manner before output generation. This is especially beneficial for latency reduction in long-context tasks, e.g., document summarization. For RNNs, this overlap applies at the first time step of inference. Moreover, while SLIDE focuses on full model downloads, it can extend to cases where users update models by downloading task-specific parameters (e.g., LoRA matrices for each LLM layer [9]).
@@ III-C2 Layer inference latency
where $V_{1}\left(k,l_{i}\right)$ denotes the time required for user $k$ to transfer layer $l_{i}$ from system memory to GPU memory [^2] Layer- or block-wise parameter transfer and loading is supported by modern deep learning frameworks, e.g., PyTorch [32]. Without loss of generality, the GPU memory transfer time is represented by $V_{1}(k,l_{i})$, typically approximated as the ratio of the data size of layer $l_{i}$ to the data rate between the system memory and GPU memory [33].. The second term in (4) represents the time required for user $k$ to perform the forward propagation of layer $l_{i}$ on the GPU [34, 35, 36]. Moreover, $b_{k}$ denotes the inference batch size of the input data samples of user $k$, $W_{l_{i}}$ represents the computation workload (in FLOPs) required by layer $l_{i}$ to process a single data sample, $\kappa_{k}$ denotes the computing intensity (in cycles/FLOP) of user $k$, and $f_{k}$ is the GPU clock frequency (in cycles/s) of user $k$. $z_{k,l_{i}}\in\left[0,1\right]$ is the GPU frequency allocation scaling factor of user $k$ for the forward propagation of layer $l_{i}$, satisfying
@@ III-E Task Throughput Maximization Problem
In addition, although $\mathcal{P}1$ is formulated for a single decision time slot, long-term user fairness across multiple time slots can be incorporated without changing the core procedure of SLIDE. This can be achieved by introducing time-varying user weights that increase for previously unserved users [41], and by sequentially solving the per-slot optimization problem. Since the user weights are known at the beginning of each decision time slot, the per-slot optimization problem preserves the same problem property as $\mathcal{P}1$, without affecting the subsequent algorithm design.
@@ IV-A1 Computing resource allocation for user $k$ with model $i$ given $y_{k}$
IV-A2 Model provisioning and computing resource allocation for user $k$ given $y_{k}$
@@ V-A Experimental Setup
  • Conventional DAI: the conventional downloading and inference approach, where users start inference only after the entire model is downloaded [19]. To ensure a fair comparison, this baseline adopts the same user selection, bandwidth allocation, and computing resource allocation procedures as SLIDE by solving problem $\mathcal{P}1$ using Algorithm 2, with the only difference lying in the E2E latency model. Specifically, according to (6), the E2E latency of user $k$ in conventional DAI is given by $t_{k,L_{i}}\left(y_{k},{\bf{Z}}_{k,i}\right)=\sum\limits_{l_{i}=1}^{L_{i}}\frac{S_{l_{i}}}{y_{k}BR_{k}}+\sum\limits_{l_{i}=1}^{L_{i}}T_{k,i}\left(z_{k,l_{i}}\right)$, which is the sum of model downloading and inference latencies, corresponding to the sequential model downloading and inference.